OpenStudy (anonymous):

b+9/b-7 divided by 6b+54/b-1

jhonyy9 (jhonyy9):

b+9/b-7 / 6b+54/b-1 =b+9/b-7 x b-1/6(b+9) =b-1/(b-7)6